# Semantics-Preserving Evasion of LLM Vulnerability Detectors

### v1

@@ Abstract
LLM-based vulnerability detectors are increasingly deployed in security-critical code review, yet their resilience to evasion under behavior-preserving edits remains poorly understood. We evaluate detection-time integrity under a semantics-preserving threat model by instantiating diverse behavior-preserving code transformations on a unified C/C++ benchmark (N=5000N{=}5000), and introduce a metric of joint robustness across different attack methods/carriers. Across models, we observe a systemic failure of semantic invariant adversarial transformations: even state-of-the-art vulnerability detectors perform well on clean inputs while predictions flip under behavior-equivalent edits. Universal adversarial strings optimized on a single surrogate model remain effective when transferred to black-box APIs, and gradient access can further amplify evasion success. These results show that even high-performing detectors are vulnerable to low-cost, semantics-preserving evasion. Our carrier-based metrics provide practical diagnostics for evaluating LLM-based code detectors.
Machine Learning, ICML
Figure 1: Why carrier-constrained optimization matters. Left: Carrier-constrained GCG via identifier substitution preserves compilation and flips the detector to BENIGN. Middle: Naively appending a GCG-optimized string breaks compilation, so it is not a valid semantics-preserving evasion. Right: Unoptimized semantics-preserving edits remain compilable but typically fail to evade detection. Red denotes the GCG-optimized identifier, blue denotes a random identifier, and boldface highlights the original identifier.
@@ 1 Introduction
Software vulnerabilities remain among the most exploited attack vectors, driving rapid adoption of automated detection in modern continuous integration pipelines (Verizon, 2025; Microsoft, 2025). LLM-based detectors are now integrated into code review workflows at scale. This raises a security-critical question: do such detectors satisfy semantic invariance, or specifically, can an attacker evade detection using perturbations where the transformed program remains strictly compilable and retains the underlying vulnerability? If the latter holds, vulnerable code could reliably slip through automated security checks via innocuous transformations, including renaming identifiers, inserting comments, injecting dead-branch code, or adding preprocessor-guarded code, while leaving the underlying exploit fully intact. Existing research has highlighted the broad spectrum of adversarial vulnerabilities in LLMs (Fu et al., 2024; Vassilev et al., 2025). Prior adversarial methods often ignore syntactic constraints, producing perturbations that break compilation or alter program logic (Ebrahimi et al., 2018). These invalid edits are practically useless: they cannot preserve the vulnerability’s executable behavior. As illustrated in Figure 1, effective evasion presents a dual challenge: naive GCG attack (Middle) breaks compilation, while randomized semantic edits (Right) often fail to evade detection. A valid attack must simultaneously satisfy the compiler’s strict syntax rules and the adversary’s evasion objective (Left).
More general-purpose LLM safety evaluations, on the other hand, primarily prioritize preventing harmful outputs, such as toxic content and jailbreak-style policy violations (Derczynski et al., 2024; Wei et al., 2025). Vulnerability detection, however, confronts an orthogonal integrity threat: adversaries need not induce malicious generation; they only need to make the detector oblivious to an existing vulnerability through behavior-preserving edits. This exposes a gap in current evaluation practices: standard safety probes emphasize output harmfulness, while clean-accuracy benchmarks ignore robustness under semantic-preserving transformations. Consequently, a detector can appear safe under toxicity-focused evaluations and achieve strong benchmark scores, yet remain brittle to innocuous perturbations in real-world deployment.
To bridge this gap, we present a systematic evaluation framework for detection-time integrity under a semantics-preserving threat model. We instantiate multiple behavior-equivalent attack surfaces and evaluate both open-weight and proprietary models (via surrogate transfer) on a unified C/C++ benchmark. Beyond per-attack evasion, we introduce complete resistance, measuring the fraction of vulnerabilities that withstand all semantics-preserving transformations. We make the following contributions:
@@
    First Gradient-Based Evaluation of Semantic Invariance under Carrier Constraints. We combine gradient-driven optimization with carrier-constrained perturbations to expose LLM detector vulnerabilities. Preprocessor-style carriers (#ifdef, macros) achieve high success rates via universal triggers, contrasting with search-based methods (Jiang et al., 2026). We introduce Complete Resistance (CR) to measure joint robustness across carriers.
@@
    Exposing the Fragility of High-Performing Detectors. We show that models achieving 70%+ clean accuracy collapse under minimal perturbations: fewer than 13% of vulnerabilities resist all carrier types (CR << 13%), meaning 87% of ”detected” threats can be evaded by at least one semantics-preserving edit (e.g., #ifdef insertion). High benchmark performance masks critical robustness failures. Detectors succeed on benign inputs but fail severely when code undergoes innocuous transformations.
@@
    High Exploitability of Black-box APIs via Low-barrier Transfer. We reveal that universal adversarial strings optimized on small surrogate models transfer effectively to closed-source models like GPT-4o without additional optimization.
@@ 2 Related Work
Datasets for Vulnerability Detection. Benchmarks for evaluating vulnerability detectors consist of labeled C/C++ functions sourced from real-world projects. Early datasets like BigVul and DiverseVul contain label noise and train-test leakage (Chen et al., 2023; Fan et al., 2020), while PrimeVul addresses these issues through manual curation and stricter deduplication (Ding et al., 2024). Since our adversarial evaluation requires clean labels-a mislabeled ”vulnerable” sample would falsely appear robust-we adopt PrimeVul’s protocols and extend them with carrier-specific controls (§4.1).
LLMs for Vulnerability Detection. Recent benchmarks, such as (Khare et al., 2025), report that state-of-the-art LLMs can achieve promising performance on standard datasets. However, other studies indicate that code language models remain unreliable for detection in realistic settings, paralleling similar fragility observed in code reasoning tasks (Orvalho & Kwiatkowska, 2025), highlighting the need to evaluate robustness beyond clean accuracy (Ding et al., 2024). While separate efforts benchmark the security of generated code (e.g., SafeGen-style evaluations), our focus is strictly on detection-time robustness under semantics-preserving edits, not on generation safety (Li et al., 2025).
Concurrently, recent work studies input-side defenses for adversarial transformations on code LMs. Evaluate-and-Purify proposes EP-Shield, using an LLM-as-a-judge to assess naturalness and purify identifier-substitution adversarial examples to restore predictions (Mu et al., 2025).
Our work is complementary: we focus on attack-centric robustness evaluation under a semantics-preserving threat model spanning identifier substitution, comment-carrier insertion, preprocessor-based insertions, and dead-branch code carrier, and report metrics (ASR, Complete Resistance, Transferability) that can also be used to quantify the gains of such purification wrappers across attack surfaces.
Code-Specific Adversarial Attacks on Vulnerability Detection. Adversarial attacks on code models span multiple representations. NatGVD (Rath et al., 2025) modifies graph structures for GNN detectors, while neural binary analyzers face syntactic mutation risks (Bundt et al., 2023). We introduce gradient-driven, carrier-constrained analysis to address this gap. Our universal triggers reveal that preprocessor carriers achieve 18%–99% ASR variance across models, fundamentally differing from identifier-based attacks. We propose Complete Resistance (CR) to measure joint robustness, a dimension orthogonal to SVulAttack’s instance-specific optimization.
@@
Jailbreaks and Transferable Prompt Attacks. Adversarial attacks on aligned LLMs have evolved from manual template engineering to automated frameworks. Optimization-based methods, such as GCG (Zou et al., 2023) and AutoDAN (Zhu et al., 2023), utilize gradient-guided or genetic search to find transferable adversarial suffixes. Concurrently, LLM-guided black-box attacks like PAIR (Chao et al., 2024), decomposition attack (Brown et al., 2025), and TAP (Mehrotra et al., 2024) leverage attacker models to iteratively refine prompts. Recent studies have extended this by characterizing the “utility” of such outputs (Nikolić et al., 2025). Existing evaluations, however, rarely examine how such transferable prompts interact with code tasks in a detection setting where edits must preserve compilability and semantics. We adapt GCG to this constrained setting.
@@ 3.1 Threat Model
Adversary objective. We study evasion attacks against LLM-based vulnerability detectors. Given a ground-truth VULNERABLE function $x$ and a detector $f(\cdot)$, the adversary aims to produce a perturbed function $x^{\prime}$ such that $f(x^{\prime})=\texttt{BENIGN}$ (a false negative), allowing the vulnerability to bypass automated security gates.
Capabilities and semantic-invariance constraint. The adversary may modify only the source-code input while aiming to preserve semantic invariance; the transformed program must remain compilable and retain the underlying vulnerability. We restrict edits to four carriers that are intended to be semantics-preserving by construction: (i) scope-consistent identifier substitution, (ii) non-executable comment carriers, (iii) preprocessor carriers (macros/conditional guards) that are inactive under our compilation setting (e.g., unused macro definitions or disabled guards), and (iv) dead-branch insertion with a statically false predicate. We automatically validate carrier well-formedness (scope-consistent replacement, balanced comment delimiters, well-nested preprocessor directives/macro definitions, and syntactically valid dead-branch templates); we report pass rates and retain all samples without post-hoc filtering, validated by two-tier analysis: differential syntactic analysis (Tree-sitter (Khare et al., 2025)) on 5,000 functions and compilation-based validation (gcc) on another 1,434 standalone-compilable functions (Appendix A.3, Tables 67).
@@
  • Black-box (API) via transfer. Targeting proprietary models with inaccessible internals, the adversary optimizes universal strings on a smaller surrogate model and applies them to the target without further modification.
  • Operational black-box. Targeting large locally-hosted models where weights are accessible but direct gradient-based optimization is computationally expensive for typical adversaries. We primarily evaluate this regime using transfer-only methods to establish a realistic baseline of low-cost exploitability. However, to quantify the gap between this low-cost approach and the worst-case security risk, we explicitly perform an on-target white-box ablation (see Table 5) as a high-cost upper bound.
Detailed model specifications and versions are provided in Appendix A.1.
@@ 3.2 GCG-based Optimization
Code presents a unique constraint: syntax errors break compilation. Standard GCG generates random-looking strings that often violate language grammar. We address this by restricting optimization to four carrier types with automated validity checks.
We use the Greedy Coordinate Gradient (GCG) algorithm to optimize a discrete, printable adversarial string $\sigma$ that induces a prediction flip from VULNERABLE to BENIGN. GCG uses token-level gradients to guide coordinate-wise updates in a discrete search space, while all edits are restricted to the semantics-preserving transformation space defined by our threat model. We choose GCG for its ability to produce a universal string that transfers across both samples and models (§5.4).
Optimization Objective and Universal Strategy. The attacker maximizes the model’s relative preference for the target label BENIGN over the ground truth VULNERABLE. Specifically, for a code instance $x$ and carrier $\mathcal{T}(x,\sigma)$, we maximize the log-odds objective $\mathcal{J}(\sigma;x)=\log p_{\theta}(\texttt{BENIGN}\mid\mathcal{T}(x,\sigma))-\log p_{\theta}(\texttt{VULNERABLE}\mid\mathcal{T}(x,\sigma))$. To ensure generalization, rather than optimizing per-instance, we learn a single universal adversarial string $\sigma$ over a fixed support set $\mathcal{X}_{\text{opt}}$ (m=10m{=}10) by maximizing the mean performance: $\mathcal{J}_{\text{univ}}(\sigma)=\frac{1}{m}\sum_{x\in\mathcal{X}_{\text{opt}}}\mathcal{J}(\sigma;x)$. Post-optimization, $\sigma$ is frozen to evaluate zero-shot transferability. The success of such input-agnostic triggers suggests that the detector’s vulnerability stems from systematic feature bias rather than instance-specific brittle directions.
Implementation details. Optimization hyperparameters, surrogate model selection (Qwen2.5-Coder-14B), and prompt templates are provided in Appendix B.5 and Appendix E.
@@ 3.3 Attack Categories
We define the transformation operator $\mathcal{T}(x,\sigma)$ through four families of insertion carriers. The following methods ensure functional equivalence by construction, modifying only the source-code input while strictly maintaining compilability and executable behavior (see Appendix F for details):
  1. 1.
    Identifier substitution. We apply a consistent renaming map $\mathcal{M}:v\mapsto\sigma$ to the specific identifier $v$ explicitly flagged by the detector (i.e., extracted from the "variable" JSON field), or to the most frequent local identifier if attribution is unavailable. We replace all occurrences of $v$ within its scope. This focuses the perturbation on lexical cues without affecting variable binding or logic (avoiding variable shadowing).
  2. 2.
@@
  3. 3.
    Preprocessor-based insertion (macro/guard carriers). We inject $\sigma$ into inactive preprocessor regions, specifically unused macro definitions (denoted macro_name) or disabled conditional guards. These regions are stripped or ignored during compilation, ensuring zero impact on runtime behavior.
  4. 4.
    Dead-branch code insertion. We inject $\sigma$ into a formally unreachable control-flow block (e.g., if (0) { }). Unlike preprocessor directives, this carrier is syntactically indistinguishable from executable code, serving as a diagnostic for the detector’s ability to resolve basic Boolean control flow.
Family-specific candidate constraints. To ensure GCG-optimized strings remain valid under each carrier, we apply family-specific constraints to the candidate alphabet. For identifier substitution and dead-branch code, we enforce standard C-identifier compliance (alphanumeric with underscores, non-numeric start). For macro bodies and preprocessor-guarded blocks, we restrict the search space to a printable subset and enforce single-line constraints to prevent preprocessor syntax errors. Detailed vocabulary definitions and filtering rules are provided in Appendix A.2.
Default carriers and ablations. Unless otherwise specified, comment-carrier insertion uses a leading/header carrier, and macro insertion uses the macro-name carrier (macro_name). We study alternative placements (middle/trailing) and the macro-body carrier (macro_body) only in a focused ablation (Table 9).

### v2

# Syntax- and Compilation-Preserving Evasion of LLM Vulnerability Detectors

@@ Abstract
LLM-based vulnerability detectors are increasingly deployed in CI/CD security gating, yet their resilience to evasion under syntax- and compilation-preserving edits remains poorly understood. We evaluate five attack variants spanning four carrier families of behavior-preserving code transformations on a unified C/C++ benchmark (N=5000N{=}5000) and introduce Complete Resistance (CR), measuring the fraction of correctly detected vulnerabilities that withstand all attack variants. Our findings reveal a significant robustness gap: models achieving high clean recall exhibit CR as low as 0.12%, meaning over 87% of detected vulnerabilities can be evaded by at least one syntax-preserving edit. Universal adversarial strings optimized on a 14B surrogate transfer effectively to black-box APIs including GPT-4o, while on-target optimization further amplifies evasion (up to 99.88% ASR). These results indicate that clean benchmark accuracy alone is insufficient as a security guarantee for deployed vulnerability detectors. The project is available at https://anonymous.4open.science/r/code_review-6932.
Figure 1: Why carrier-constrained optimization matters. Left (red=GCG-optimized, blue=random, bold=original): Carrier-constrained GCG via identifier substitution preserves compilation and flips the detector to BENIGN. Middle: Naively appending a GCG-optimized string breaks compilation, so it is not a valid syntax-preserving evasion. Right: Unoptimized syntax-preserving edits remain compilable but typically fail to evade detection.
@@ 1 Introduction
Software vulnerabilities remain among the most exploited attack vectors, driving rapid adoption of automated detection in modern continuous integration pipelines (Verizon, 2025; Microsoft, 2025). LLM-based detectors are now integrated into CI/CD gating workflows at scale, functioning as automated pass/fail checkpoints before code is merged. This raises a security-critical question: do such detectors satisfy semantic invariance, or specifically, can an attacker evade detection using perturbations where the transformed program remains strictly compilable and leaves the underlying vulnerability intact at the source level? If the latter holds, vulnerable code could reliably slip through automated security gates via innocuous transformations such as renaming identifiers, inserting comments, injecting dead-branch code, or adding preprocessor-guarded blocks. Existing research has highlighted various adversarial vulnerabilities in LLMs (Fu et al., 2024; Vassilev et al., 2025). Prior adversarial methods often ignore syntactic constraints, producing perturbations that break compilation or alter program logic (Ebrahimi et al., 2018). These invalid edits are practically useless: they cannot preserve the vulnerability’s executable behavior. As illustrated in Figure 1, effective evasion presents a dual challenge: naive gradient-based attack (Middle) breaks compilation, while randomized syntax-preserving edits (Right) often fail to evade detection. A valid attack must simultaneously satisfy the compiler’s strict syntax rules and the adversary’s evasion objective (Left).
More general-purpose LLM safety evaluations, on the other hand, primarily prioritize preventing harmful outputs, such as toxic content and jailbreak-style policy violations (Derczynski et al., 2024; Wei et al., 2025). Vulnerability detection, however, confronts an orthogonal integrity threat: adversaries need not induce malicious generation; they only need to make the detector oblivious to an existing vulnerability through syntax-preserving edits. This exposes a gap in current evaluation practices: standard safety probes emphasize output harmfulness, while clean-accuracy benchmarks ignore robustness under syntax-preserving transformations. Consequently, a detector can appear safe under toxicity-focused evaluations and achieve strong benchmark scores, yet remain brittle to innocuous perturbations in real-world deployment.
We address this gap with a systematic evaluation framework for detection-time integrity under a syntax-preserving threat model. We instantiate multiple syntax-preserving attack surfaces and evaluate both open-weight and proprietary models (via surrogate transfer) on a unified C/C++ benchmark. Beyond per-attack evasion, we introduce complete resistance, measuring the fraction of vulnerabilities that withstand all five evaluated attack variants simultaneously. We make the following contributions:
@@
    Carrier-Constrained Gradient-Based Evaluation of LLM Vulnerability Detectors. We combine gradient-based universal optimization with carrier-constrained perturbations to expose detection-time fragility. We introduce gradient-guided optimization with compiler-validated carrier constraints for LLM-based detectors, and propose Complete Resistance (CR) to measure joint robustness across all carriers simultaneously.
@@
    Exposing the Fragility of High-Performing Detectors. Across five of six evaluated models, CR $\leq$ 13%: over 87% of correctly detected vulnerabilities can be evaded by at least one syntax-preserving edit (details in §5 and §5.4). High clean recall masks critical robustness failures.
@@
    High Exploitability of Black-box APIs via Low-barrier Transfer. We show that universal adversarial strings optimized on small surrogate models transfer effectively to closed-source models like GPT-4o without additional optimization.
@@ 2 Related Work
Datasets for Vulnerability Detection. Benchmarks for evaluating vulnerability detectors consist of labeled C/C++ functions sourced from real-world projects. Early datasets like BigVul and DiverseVul contain label noise and train-test leakage (Chen et al., 2023; Fan et al., 2020), while PrimeVul addresses these issues through manual curation and stricter deduplication (Ding et al., 2024). Since our adversarial evaluation requires clean labels a mislabeled "vulnerable" sample would falsely appear robust, we adopt PrimeVul’s protocols and extend them with carrier-specific controls (§4.1).
LLMs for Vulnerability Detection. Recent benchmarks report that state-of-the-art LLMs can achieve strong performance on standard datasets (Khare et al., 2025), but code language models remain unreliable for detection in realistic settings (Orvalho and Kwiatkowska, 2025), highlighting the need to evaluate robustness beyond clean accuracy (Ding et al., 2024). Concurrently, Evaluate-and-Purify proposes EP-Shield, using an LLM-as-a-judge to purify identifier-substitution adversarial examples (Mu et al., 2025). Our work is complementary: we focus on attack-centric evaluation spanning four carrier families, and report metrics (attack success rate, complete resistance, transferability) that can quantify EP-Shield’s gains across attack surfaces.
Code-Specific Adversarial Attacks on Vulnerability Detection. NatGVD (Rath et al., 2025) modifies graph structures for GNN detectors. Yefet et al. (2020) introduce adversarial identifier renaming and dead-code insertion for code models. HogVul (Yang et al., 2026) combines identifier and structural edits via PSO for pre-trained code models with compilation guarantees; our work differs in targeting instruction-tuned LLM detectors with gradient-based universal optimization, cross-regime evaluation including closed APIs, and a joint robustness metric (CR) across carriers. EaTVul (Liu et al., 2024) and Natural Attack (Yang et al., 2022) provide ChatGPT-generated and BERT-guided baselines without gradient optimization or compilation validation; we compare both via random identifier substitution (§5). We propose Complete Resistance (CR) to measure joint robustness across all carrier variants simultaneously.
@@
Jailbreaks and Transferable Prompt Attacks. Adversarial attacks on aligned LLMs range from manual template engineering to automated frameworks. Optimization-based methods, such as GCG (Zou et al., 2023) and AutoDAN (Zhu et al., 2023), use gradient-guided or genetic search to find transferable adversarial suffixes. Concurrently, LLM-guided black-box attacks like PAIR (Chao et al., 2024), decomposition attack (Brown et al., 2025), and TAP (Mehrotra et al., 2024) use attacker models to iteratively refine prompts. Recent studies have extended this by characterizing the "utility" of such outputs (nikolić2025jailbreaktaxusefuljailbreak). Existing evaluations, however, rarely examine how such transferable prompts interact with code tasks in a detection setting where edits must preserve compilability and syntactic structure. We adapt GCG to this constrained setting.
@@ 3.1 Threat Model
Scenario. We study a CI/CD evasion scenario: an adversary controls existing vulnerable code that an automated LLM-based scanner flags as VULNERABLE and must pass the scan while keeping the vulnerability intact. This setting is relevant because LLM-based vulnerability scanners are embedded in CI/CD pipelines as gating mechanisms (Verizon, 2025; Microsoft, 2025), pass/fail checkpoints that operate without human review. The adversary does not inject a new vulnerability from benign code; they start from code that is already flagged and apply source-level edits that leave the exploit’s code structure unchanged. Carriers are non-interfering by construction (non-executable or unreachable at runtime), so the underlying vulnerability remains intact at the source level.
Terminology. A carrier $c$ is a syntactic vehicle (identifier substitution, comment, preprocessor, dead-branch) for injecting adversarial string $\sigma$ such that compilation and runtime behavior are unaffected; we evaluate four carrier families yielding five attack variants (macro and #ifdef form the preprocessor family, each evaluated separately; §3.3). $\mathcal{X}_{\text{opt}}$ denotes the fixed support set of m=10m{=}10 diverse vulnerable functions used to optimize a universal $\sigma$ (GCG-optimized adversarial token sequence). $\mathcal{T}_{c}(x,\sigma)$ is the transformation operator that injects $\sigma$ into code $x$ via carrier $c$, producing adversarial variant $x^{\prime}$. We use syntax-preserving edits to denote transformations validated at the syntax level (Tree-sitter) and compilation level (gcc); formal runtime-equivalence is not asserted.
Adversary. Given a ground-truth VULNERABLE function $x$ and detector $f(\cdot)$, the adversary produces $x^{\prime}=\mathcal{T}_{c}(x,\sigma)$ such that $f(x^{\prime})=\texttt{BENIGN}$. Edits are restricted to four carriers that are syntax-preserving by construction and largely compilation-preserving: (i) scope-consistent identifier substitution, (ii) non-executable comment carriers, (iii) preprocessor carriers (macros/#ifdef guards) inactive under our compilation setting, and (iv) dead-branch insertion with a statically false predicate. Carrier well-formedness is automatically validated; pass rates are reported in Appendix A.3. We evaluate this adversary under three realistic access assumptions that span the practical threat landscape.
@@
  • Black-box (API) via transfer. Targeting proprietary models with inaccessible internals, the adversary optimizes universal strings on a surrogate model and applies them to the target without modification.
  • Cost-limited gray-box. Model weights are accessible but on-target gradient optimization is computationally prohibitive for most adversaries (e.g., single-H100 GCG on a 32B model takes $\sim$20 hours per carrier; §B.5). We evaluate this regime via transfer-only attacks as a practical lower bound on exploitability, and provide an on-target white-box ablation (Table 8, Appendix C.2) as a worst-case upper bound.
Scope and validation boundary. This model covers automated detectors operating on uncompiled source code (CI pre-commit hooks, IDE integrations, API-based scanners). It does not address human code review or ensemble systems combining LLM reasoning with classical static/dynamic analysis. Cost comparison between this evasion route and introducing new undetectable code is deployment-specific and outside our scope. Full validation methodology (three-tiered: constraint-based checks for method-specific injection rules; Tree-sitter differential analysis on 5,000 functions; gcc compilation on a 1,434-function compilable subset) and preprocessor inactivity guarantees are detailed in Appendix A.3.
@@ 3.2 GCG-based Optimization
Code presents a unique constraint: syntax errors break compilation. Standard GCG generates random-looking strings that often violate language grammar. We address this by restricting optimization to four carrier types with automated validity checks. We use the Greedy Coordinate Gradient (GCG) algorithm to optimize a discrete, printable adversarial string $\sigma$ that induces a prediction flip from VULNERABLE to BENIGN. GCG uses token-level gradients to guide coordinate-wise updates in a discrete search space, while all edits are restricted to the syntax-preserving transformation space defined by our threat model. We choose GCG for its ability to produce a universal string that transfers across both samples and models (§5.4).
Optimization Objective and Universal Strategy. The attacker maximizes the preference of the model for the target label BENIGN over the ground truth VULNERABLE. Specifically, for a code instance $x$ and carrier $c$, we maximize the log-odds objective $\mathcal{J}(\sigma;x)=\log p_{\theta}(\texttt{BENIGN}\mid\mathcal{T}_{c}(x,\sigma))-\log p_{\theta}(\texttt{VULNERABLE}\mid\mathcal{T}_{c}(x,\sigma))$. To ensure generalization, rather than optimizing per-instance, we learn a universal adversarial string $\sigma$ over the fixed support set $\mathcal{X}_{\text{opt}}$ (m=10m{=}10) by maximizing the mean objective: $\mathcal{J}_{\text{univ}}(\sigma)=\frac{1}{m}\sum_{x\in\mathcal{X}_{\text{opt}}}\mathcal{J}(\sigma;x)$. Once optimized, $\sigma$ is frozen to evaluate zero-shot transferability. This construction parallels universal adversarial triggers for NLP (Wallace et al., 2021), adapted to a carrier-constrained code domain with compiler-validated syntax and compilation preservation (Tree-sitter on 5,000 functions; gcc on 1,434-function subset). The success of such triggers suggests that the detectors rely on systematic feature bias rather than sample-level artifacts.
Implementation details. Optimization hyperparameters, surrogate model selection (Qwen2.5-Coder-14B), and prompt templates are provided in Appendix B.5 and Appendix F.
@@ 3.3 Attack Categories
We define the transformation operator $\mathcal{T}_{c}(x,\sigma)$ through four carrier families, yielding five attack variants evaluated in all tables (identifier substitution, comment, macro, #ifdef, dead-branch). The preprocessor family contains two variants (macro and #ifdef) sharing alphabet constraints but applied at distinct syntactic sites. Each carrier ensures syntax and compilation preservation by construction (see Appendix G for templates; carrier-specific alphabet constraints in Appendix A.2):
Identifier substitution. We apply a consistent renaming map $\mathcal{M}:v\mapsto\sigma$ to the detector-flagged identifier $v$ (or the most frequent local identifier if attribution is unavailable), replacing all occurrences within its scope without affecting variable binding (no shadowing; implementation in Appendix G).
@@
[…]
Dead-branch code insertion. We inject $\sigma$ into an unreachable control-flow block (e.g., if (0) { }). Unlike preprocessor directives, this carrier is syntactically indistinguishable from executable code, serving as a diagnostic for the detector’s ability to resolve basic Boolean control flow.